Last year Charlie Investments issued a 10 year, 12% semiannual coupon bond at it's par value of $1000. Currently the bond can be paid in 4 years at a price of $1060 and it sells for $1100. Calculate the bond's nominal yield to maturity and it's nominal yield to call. Would an investor be more likely to earn the YTM or the YTC? Briefly explain your answer.

1
Expert's answer
2022-03-14T13:24:40-0400

Solution


Coupon Bond's Par/ Face Value "=\$1,000"


Annual Rate "=12\%"

Semiannual rate "=\frac{12\%}{2}=6\%"


Maturity Period "=10" years


Years to call "= 4" yrears


Call-able Price "=\$1060"


Call-able Sale Price "=\$1100"



We know the formula,


Annual Coupon Payment "=" Par value "\times" Annual coupon rate


Replacing vales, we get


Annual Coupon Payment "=\$1000\times 12\%"


"=\$120"


Now, the current yield is "=\frac{annual\ coupon\ payment}{current\ price}"



"=\frac{\$120}{\$1100}\\=0.1090909....\\
=10.91\%"




Bond's nominal yield to maturity is calculated by taking promised interest rate and multiplying by the number of years until maturity


Therefore,


Bond's nominal yield to maturity "=12\%\times 10 = \$ 120"



Yield to Call is calculated as


"= \frac{{Annual\,Coupon\,Yield + \frac{{\left( {Call\,price - Bond\,face\,value} \right)}}{{Number\,of\,years\,to\,call}}}}{{\frac{{\left( {Call\,price + Bond\,Bond\,face\,value} \right)}}{2}}}"


"= \frac{{120 + \frac{{\left( {1060 - 1100} \right)}}{4}}}{{\frac{{\left( {1060 + 1100} \right)}}{2}}} = \frac{{120 - 10}}{{1080}} = 0.10185185185...\\
 = 10.85\%"